OpenStudy (anonymous):

Teresa and Julia are among 10 students who have applied for a trip to Washington DC. Two students from the group will be selected at random for the trip. What is the probability that Teresa and Julia will be the two students selected?

OpenStudy (anonymous):

Probability = number of specific cases / all possibilities

OpenStudy (anonymous):

How many total number of possibilites for 2 students are there? @mayanmystery22

OpenStudy (anonymous):

what you're saying is exactly what I did. It's initially 1 out of 10 students. Then after one was chosen, it became 1 out of 9 students to pick the second one. I would think it would be 1/90 since I multiply the two together, but that is not one of my answer choices.

Directrix (directrix):

Order of selection does not matter. So, the probability that T will be chosen is 1/10 followed by a probability that J will also be chosen of 1/9. That product is 1/90. But, J could be chosen first and then T which also yields a probability of 1/90. P ( T and J) = 1/90 + 1/90 = 2/90 = ? --> Is this an option?

Directrix (directrix):

I looked at combinations. C(2,2) means the ways to select the T and J where order of selection does not count. C(8,0) means the ways to select zero of the other 8 students. C(2,2) * C(8,0) is divided by C(10,2) which is the number of ways to select two people for the trip. That also cranks out to the same result as above. @mayanmystery22

Directrix (directrix):

Check out #65 at the following link: http://tinyurl.com/cf7yzzg